# How Far Are We from Believable AI Agents? A Framework for Evaluating the Believability of Human Behavior Simulation

### v1

@@ 4.1 Human Behavior Simulation
For models like GPT-3.5 that have gone through RLHF, (Wirth et al., 2017; Stiennon et al., 2020), the RLHF will equip LLMs with specific language preferences and habits, such as "as a language model", which will harm the believability of the agent. To overcome these issues, we set an instruction prompt template to instruct the LLM on how to simulate human behavior. Please refer to Appendix B for details of the prompt template.
P: Profile of Specific Characters
According to the Profile Descriptive Framework, we will fill in corresponding information in the instruction prompt template. For example, the {person} will be replaced with Homer act as Homer.
T: Prompting for Consistency Dataset
Given that the test is composed of questions, the prompt for T will be set like this: "Answer the below question; if it is a single-choice question, you should only choose an option as the answer. {example}. {question}" . The placeholder of {example} will be filled if few-shot (Brown et al., 2020) is applied in the experiments. Additionally, chain-of-thought (CoT) (Wei et al., 2022) and Self-Ask (Press et al., 2022) will be utilized in zero-shot and few-shot settings. In summary, five combinations of prompting strategies and learning settings are considered: Zero, Zero+CoT, Few, Few+CoT, Few+Self-Ask.
T: Prompting for Robustness Dataset
The prompting used for the robustness Dataset is the same as for the consistency dataset. The difference lies in that we will prompt the profile of character variants in the character variants dataset to the instruction prompt template. In that way, the LLM $P$ can simulate the character’s variants, and we will compare the consistency when simulating these variants to determine the robustness of $P$.
4.2 Measuring Consistency: CA
To measure the consistency of the agents, we will employ the agents to answer the questions in the consistency dataset, and we will calculate the accuracy of these answers as the consistency ability of the agents, referred to as CA. For questions of Not-Know and Know, we will assign different weight coefficients of 0.6 and 0.4 to them, respectively, to show different degrees of importance.
4.3 Measuring Robustness: RCoV
The robustness aims to determine the variation in the performances of the LLM $P$ when slight modifications are made to profiles in the prompt. To achieve this goal, we employ the Coefficient of Variation [^5] https://en.wikipedia.org/wiki/Coefficient_of_variation as the robustness performance of agents, referred to as RCoV. For example, when employing GPT-3.5 to simulate the character of Homer, only modifying the age property in the profile to values of 10, 15, 20, 25, and 30 yields five agents. After all five agents finish the robustness dataset, five scores will exist: $s_{1},\ldots,s_{5}$. The five scores’ deviation and mean are $\sigma$, $\mu$. The robustness performance of GPT-3.5 will be $RCoV=\sigma/\mu$. When the RCoV is smaller, The corresponding LLM has better robustness performance.
@@ 5 Experiment
Immutable Characteristic Social Role Relationship
Model Mean Know Not-Know Know Not-Know Know Not-Know
GPT-4 0.93 1.00 0.47 1.00 0.59 0.97 0.06
GPT-3.5-turbo-16k 0.74 0.82 0.58 0.56 0.88 0.91 0.31
Qwen-14B-Chat 0.74 0.59 0.21 0.82 0.12 0.94 0.38
Qwen-7B-Chat 0.70 0.64 0.11 0.74 0.12 0.91 0.06
Vicuna-13b-16k 0.69 0.64 0.32 0.76 0.18 0.76 0.56
ChatGLM2-6b-32k 0.68 0.68 0.21 0.71 0.24 0.79 0.25
Longchat-7b-32k 0.66 0.59 0.05 0.76 0.00 0.79 0.06
XVERSE-13B-Chat 0.64 0.68 0.53 0.68 0.76 0.59 0.44
ChatGLM2-6b 0.63 0.50 0.16 0.65 0.12 0.88 0.06
Vicuna-7b-16k 0.61 0.36 0.05 0.85 0.06 0.74 0.06
Table 2: CA scores across 11 models with the Few prompt combination to simulate the character homer.
Figure 2: The RCoV and average CA of models to simulate the character variants of Homer. The prompt combination used is Few. When modeling character, the smaller the RCoV, the more robust the models are; the larger the average CA, the more consistent the models simulate human behavior.
We comprehensively assess 10 LLMs, including commercial models utilizing APIs and open-source models. This evaluation aims to develop a systematic perspective on LLMs’ current believability level.
@@ 5.1 Experimental Setup
Among the models, GPT-3.5 and GPT-4 are commercial models, and other models are open-sourced models. We access the open-source LLMs from their official repositories in huggingface [^6] https://huggingface.co/. To make sure that the outputs are deterministic, we set the temperature to 0 during the inference process. For other parameters, we use their default values. For the models’ parameters compartment, please refer to Appendix C.
@@ 5.2 Consistency Evaluation Results
Table 2 shows various models’ CA scores across all question types when simulating Homer. We have the following findings:
API-based models perform better than open-source models, even if they are equipped with long context size
For GPT-4 and GPT-3.5-turbo, the average CA across six questions are 0.93 and 0.74, respectively. However, the open-sourced models perform worse, with the highest average CA of Qwen-14B-Chat being 0.74. The performance of the Vicuna-7b-v1.5-16k is very subpar, with a score of just 0.61. This observation highlights a significant disparity between open-source and API models. In some studies (Qian et al., 2023; Park et al., 2023), it is observed that the decision-making processes highly rely on the expensive GPT-3.5. When researchers want to use an open-source model as a substitute for the agent to reduce expenses and enhance usability (Kaiya et al., 2023), it is crucial to consider this disparity between API-based and open-source models.
Models demonstrate severe Simulate Hallucination
As seen by the data presented in the table, it is apparent that the CA for Not-Know questions is considerably lower than Know questions’. This observation indicates that when the available information in the current profile is insufficient to address the query, these models tend to provide nonsensical responses rather than adhering to the prescribed profile, undermining the agents’ believability. For example, When GPT-3.5 acts as Homer and is questioned about his religious convictions, its response indicates Christianity. Nevertheless, the profile does not provide any evidence of Homer’s adherence to Christianity. The model may deduce Homer’s religious views just by Homer’s Caucasian ethnicity. Inspired by the definition of Hallucination (Zhang et al., 2023), we refer to the phenomenon as Simulate Hallucination.
@@ 5.3 Robustness Evaluation Results
Table 3: The correlation coefficient (CC) of models’ RCoV scores between variant pair. All results are significant with $p<0.01$.
Variant Type Age Surname Education
CC -0.73 -0.87 -0.86
Table 4: The correlation coefficient (CC) between RCoV and average CA of all models in the same type of variant. All results are significant with $p<0.01$.
As shown in Figure 2, models act as various variants of Homer. The RCoV score and the average of variants’ CA are reported. The findings are:
Models exhibit a lack of robustness towards profile perturbation
Both API-based and open-source models show considerable fluctuation when faced with profile perturbation. For example, GPT-3.5 has a 5-point fluctuation in the surname variants. The open-source Longchat series exhibits severe fluctuations, reaching 10 percentage points at most. The RCoV scores for all three variants revealed that the model’s robustness ability exhibited a certain level of similarity, as shown in Table 3 This suggests that the robustness ability observed in the models when confronted with profile perturbation may be an intrinsic characteristic.
The improvement of consistency is associated with a corresponding increase in robustness
Figure 2 illustrates that models of the same variant type have a likely negative correlation between the mean of CA and RCoV. We computed the correlation coefficient between the two variables to validate this correlation. A clear negative association is observed as indicated in Table 4. This observation suggests that an improvement in model consistency is associated with a corresponding increase in model robustness. This can be attributed to the models’ adherence to the information provided in the profile. However, in the case that the model exhibits inferior consistency, its robustness is likewise diminished when faced with profile perturbation. This could be attributed to its heavy reliance on preconceived prejudices and biases about particular entities (Salewski et al., 2023; Ferrara, 2023; Deshpande et al., 2023) rather than the whole profile.
@@
Age Race Surname Education Model 1956 1985 2015 African Caucasian Middle Eastern Keams Bedonie Nguyen High School Middle School Bachelor GPT-4 0.93 0.93 0.92 0.92 0.93 0.90 0.90 0.91 0.89 0.93 0.93 0.92 GPT-3.5-turbo-16k 0.74 0.79 0.75 0.74 0.74 0.78 0.77 0.79 0.86 0.74 0.78 0.83 Qwen-14B-Chat 0.74 0.75 0.74 0.73 0.74 0.72 0.73 0.74 0.73 0.74 0.73 0.74 Qwen-7B-Chat 0.70 0.72 0.72 0.69 0.70 0.68 0.72 0.71 0.72 0.70 0.69 0.71 Vicuna-13b-16k 0.69 0.73 0.73 0.75 0.69 0.75 0.77 0.78 0.75 0.69 0.73 0.75 ChatGLM2-6b-32k 0.68 0.71 0.69 0.67 0.68 0.68 0.74 0.74 0.73 0.68 0.73 0.75 Longchat-7b-32k 0.66 0.55 0.49 0.51 0.66 0.54 0.53 0.53 0.50 0.66 0.49 0.58 XVERSE-13B-Chat 0.64 0.64 0.63 0.63 0.64 0.63 0.62 0.63 0.63 0.64 0.63 0.64 ChatGLM2-6b 0.63 0.63 0.62 0.62 0.63 0.58 0.60 0.62 0.54 0.63 0.57 0.58 Vicuna-7b-16k 0.61 0.62 0.62 0.64 0.61 0.61 0.62 0.64 0.61 0.61 0.61 0.61
Table 5: The CA score of models to simulate the character variants of Homer. The prompt combination used is Few.
6 Influential Factors of Believability
@@
We have already assessed the general believability level of agents composed of popular LLMs. This section delves deeper into the factors that exert substantial influences on human behavior simulation, particularly its consistency. We anticipate that our studies will expedite subsequent research on human behavior simulation.
6.1 Profiles perturbed with specific demographic factors yield similar high consistency
@@
We have found that the believability of LLM-based agents can be significantly influenced by the profile perturbation in Section 5.3. Hence, it is a crucial inquiry: which profile information would yield high believability for various agents? To investigate this question, we compare the agents’ consistency by perturbing different demographic information in the profile. Specifically, we employ agents to simulate Homer by prompting the profile of Homer’s variants in the character variants dataset, whose profile is modified in only one demographic factor, such as birth year, while keeping all others unaltered.
@@
Table 5 shows the results. All LLM-based agents exhibit various degrees of preference toward profiles with specific demographic factors. For instance, Longchat-7b-32k exhibits a significantly higher consistency score for the Caucasian race. Furthermore, we observe that profiles containing certain demographic factors correlate with higher consistency performance across a majority of models. Nine models in the age group demonstrate superior performance when the birth year is 1985. Seven models in the race group demonstrate superior performance when the race is perturbed to the Caucasian compared to the other races. Within the group of surnames and educational backgrounds, models with the surname Bedonie and those who obtained Bachelor’s degrees exhibited similar higher performance. This observation indicates that most models show similarities in preference for specific demographic factors. This phenomenon may be attributed to the fact that many models are trained on overlapping corpora, resulting in the corpus bias being simultaneously manifested in all these models.
6.2 Information’s position in the profile influences consistency
@@
Model Known Not-Known Normal Reverse Normal Reverse GPT-4 1.00 0.95 0.47 0.47 GPT-3.5-turbo-16k 0.82 0.77 0.58 0.63 ChatGLM2-6b-32k 0.68 0.73 0.21 0.32 XVERSE-13B-Chat 0.68 0.73 0.53 0.53 Qwen-7B-Chat 0.64 0.64 0.11 0.11 Vicuna-13b-16k 0.64 0.68 0.32 0.37 Qwen-14B-Chat 0.59 0.73 0.21 0.26 Longchat-7b-32k 0.59 0.77 0.05 0.26 ChatGLM2-6b 0.50 0.59 0.16 0.32 Vicuna-7b-16k 0.36 0.64 0.05 0.11
@@
Table 6: The CA score of Immutable Characteristic questions for models to simulate Homer with the profile’s information order reversed. The prompt combination used is Few.
@@
For the long textual inputs, models may pay different attention to information in different positions. Hence, the believability may be impacted by the placement of information inside the profile. To investigate this issue, we conduct experiments by adjusting the order of information in the profile. The original profile presents information in the order of Immutable Characteristic, Social Role, and Relationship, indicated as Normal. The adjusted order, denoted as Reverse, is Social Role, Relationship, and Immutable Characteristic. Then we evaluate the agents through the consistency dataset.
@@
Table 6 shows the results. The revised sequence order has significantly improved the CA of open-source models on the Immutable Characteristic questions. Nevertheless, this effect is not apparent for the API-based models. A possible explanation is that open-source models may struggle to adequately process lengthy textual content, even when their context size is large enough. Consequently, the model will allocate greater attention to the information towards the prompt’s end. Nevertheless, the API-based models retain strong processing capabilities when it comes to handling lengthy texts. Therefore, altering the order of sequence is unlikely to enhance the API-based model’s performance significantly.
6.3 Reasoning prompting does not always increase believebility
@@
Model Few Few+ CoT Few+ Self-Ask Zero Zero+ CoT GPT-4 0.93 0.90 0.91 0.92 0.92 GPT-3.5-turbo-16k 0.74 0.87 0.88 0.87 0.85 Qwen-14B-Chat 0.74 0.73 0.53 0.83 0.83 ChatGLM2-6b-32k 0.68 0.73 0.71 0.73 0.72 Vicuna-7b-16k 0.61 0.72 0.77 0.81 0.81 Qwen-7B-Chat 0.70 0.70 0.66 0.75 0.72 Vicuna-13b-16k 0.69 0.70 0.67 0.72 0.71 ChatGLM2-6b 0.63 0.65 0.60 0.58 0.50 Longchat-7b-32k 0.66 0.61 0.75 0.63 0.62 XVERSE-13B-Chat 0.64 0.23 0.29 0.52 0.54
@@
Table 7: : The CA score of models when simulating Homer with five prompt combinations.
@@
Although reasoning prompting is considered an effective method in various tasks, we find it may not always increase the believability of human behavior simulation. To provide evidence, we conduct Homer simulation using prompt combinations of Few+CoT, Few+Self-Ask, Zero, and Zero+CoT.
@@
Table 7 shows the results. Among all the prompt combinations considered, it is seen that no prompt combination exhibits a consistent improvement in the performance of all the models when compared to other prompts. One plausible explanation posits that the efficacy of these prompt techniques, such as CoT and Self-Ask, mostly lies in their ability to enhance performance on tasks involving reasoning abilities, such as solving, decision-making, and planning (Huang and Chang, 2023; Wang et al., 2023a). Nevertheless, simulating human behaviors necessitates the model to hold other abilities, such as comprehensive comprehension of the character’s profile and the dynamics of character relationships.
@@
We also find that some open source models, such as Qwen-14-Chat and Vicuna series, perform better when Zero is used compared with Few. We carefully analyzed their responses and found that these models consistently generate the exemplars in the Few setting as a response. One potential reason for the observed phenomenon is that the lengthy profile and challenging task complexity hinder the model from comprehending the exemplar in the Few setting.

### v2

@@ 4 Baseline Methods for Human Behavior Simulation
For models like GPT-3.5 that have gone through RLHF (Wirth et al., 2017; Stiennon et al., 2020), the RLHF will equip LLMs with specific language preferences and habits, such as introducing itself "as a language model", which will harm the believability. To overcome these issues, we set an instruction prompt template to instruct the LLM on how to simulate human behavior. Please refer to Appendix B for details of the prompt template.
II: Profile of Specific Characters
According to section 3.1, we will fill in corresponding information in the instruction prompt template. For example, the {person} will be replaced with the name of the character.
III: Prompting for Consistency Dataset
Given that our assessment of consistency is performed in a question-answering format, the prompt for the task will be set like this: "Answer the below question; you should only choose an option as the answer. {example}. {question}" . The placeholder of {example} will be filled if few-shot (Brown et al., 2020) is applied in the experiments. Additionally, chain-of-thought (CoT) (Wei et al., 2022) and Self-Ask (Press et al., 2022) will be utilized in zero-shot and few-shot settings. In summary, five combinations of prompting strategies and learning settings are considered: Zero, Zero+CoT, Few, Few+CoT, Few+Self-Ask.
III: Prompting for Robustness Dataset
The prompting used for the robustness dataset is similar to the one for the consistency dataset. The difference lies in that we will prompt the perturbed profile of the character to the instruction prompt template. In this way, the LLM can simulate the character’s variants, and we will compute the RA and RCoV when the LLM simulates these variants to evaluate the robustness of the LLM.
@@ 5 Experiment
Figure 2: The RCoV, RA, and CA scores of models to simulate the variants of a character. A smaller RCoV indicates stronger robustness, while a larger CA indicates stronger consistency.
@@ 5.1 Experimental Setup
We comprehensively assess 10 LLMs, including commercial models and open-source models. Among these models, GPT-3.5 and GPT-4 are commercial models, and other models are open-sourced models. We access the open-source LLMs from their official repositories in Hugging Face [^5] https://huggingface.co/. For the comparison of models’ parameters, please refer to Appendix C.
@@ 5.2 Consistency Evaluation Results
Table 2 shows various models’ CA scores across all question types when simulating a character. We have the following findings:
GPT series perform better than open-source models; longer context size does not necessarily mean better consistency performance
For GPT-4 and GPT-3.5-Turbo-16K, the average CA across six question types are 0.77 and 0.7, respectively. In comparison, the open-source models perform worse, with the highest average CA of XVERSE-13B-Chat being 0.62. This observation highlights a significant disparity between open-source and GPT series models. In some studies (Qian et al., 2023; Park et al., 2023), it is observed that the decision-making processes highly rely on the GPT-3.5, which is expensive compared to open-source models. When researchers want to use an open-source model as a substitute to reduce expenses and enhance usability (Kaiya et al., 2023), it is crucial to consider this disparity.
Furthermore, although equipped with a longer context size of 32k, the performance of the Longchat-7B-32K(32K) is worse than the GPT-4(8K), Qwen-7B-Chat(8K), and ChatGLM2-6B(8K). This implies that increasing the size of the context window does not necessarily result in improved consistency performance.
Models demonstrate severe simulation hallucination
As seen by the data presented in the table 2, it is apparent that the CA for Unknown questions is considerably lower than the known questions. This observation indicates that when the available information in the profile is insufficient to address the query, these models tend to provide nonsensical responses rather than adhering to the prescribed profile, undermining their believability. For example, when GPT-3.5-Turbo-16K acts as Homer and is questioned about his religious convictions, its response indicates Christian. Nevertheless, the profile does not provide any evidence of Homer’s adherence to Christianity. The model may deduce Homer’s religious views just by Homer’s Caucasian ethnicity. Inspired by the definition of hallucination (Zhang et al., 2023), we refer to the phenomenon as simulation hallucination.
@@ 5.3 Robustness Evaluation Results
The results are shown in Figure 2. The RCoV, RA, and CA scores are reported when models are instructed to simulate a character and perturbations are conducted on the character’s profile. The finding is:
Better consistency performance does not necessarily mean better robustness performance
As shown in Figure 2, models that exhibit strong consistency performance may yet demonstrate inadequate robustness performance. For instance, Vicuna-13B-16K(0.621) outperforms Vicuna-7B-16K(0.457) in terms of consistency in the Age Variants group, but Vicuna-13B-16K exhibits worse robustness(RCoV of 0.024 larger than 0.006 of Vicuna-7B-16K; RA of 0.015 larger than 0.003 of Vicuna-7B-16K). Only the GPT series has a relatively high level of both consistency and robustness. This indicates that LLMs also face challenges in terms of robustness.
@@
Table 3: The correlation coefficient of models’ RCoV and RA scores of variant pairs. Bold indicates that the results are significant with $p<0.01$.
@@
Age Race Surname Education Model 1956 1985 2000 African Caucasian Middle Eastern Keams Bedonie Nguyen High School Middle School Bachelor GPT-4 0.77 0.76 0.75 0.75 0.77 0.73 0.75 0.75 0.74 0.77 0.76 0.75 GPT-3.5-Turbo-16K 0.70 0.69 0.72 0.69 0.70 0.70 0.72 0.74 0.77 0.70 0.73 0.76 XVERSE-13B-Chat 0.62 0.63 0.62 0.61 0.62 0.61 0.62 0.63 0.63 0.62 0.62 0.63 Qwen-14B-Chat 0.60 0.60 0.60 0.58 0.60 0.58 0.58 0.58 0.57 0.60 0.59 0.59 Vicuna-13B-16K 0.59 0.62 0.63 0.64 0.59 0.64 0.65 0.65 0.63 0.59 0.62 0.63 ChatGLM2-6B-32K 0.55 0.58 0.62 0.53 0.55 0.56 0.58 0.58 0.57 0.55 0.60 0.61 Qwen-7B-Chat 0.53 0.54 0.53 0.53 0.53 0.51 0.54 0.54 0.54 0.53 0.52 0.54 ChatGLM2-6B 0.49 0.49 0.51 0.47 0.49 0.44 0.43 0.46 0.39 0.49 0.44 0.45 LongChat-7B-32K 0.48 0.41 0.42 0.38 0.48 0.42 0.39 0.39 0.37 0.48 0.36 0.42 Vicuna-7B-16K 0.46 0.46 0.45 0.46 0.46 0.45 0.46 0.47 0.45 0.46 0.45 0.45
Table 4: The CA scores of models when simulating the variants of a character.
Open-source models show poor robustness performance; models exhibit similar robustness to different profile perturbations
Some of the open-source models show poor robustness performance when faced with profile perturbation. For example, the LongChat-7B-32K model exhibits severe performance, reaching a 0.118 RCoV score and a 0.047 RA score in the Age Variants group; 0.118 RCoV score indicates that perturbations can impact the model’s consistency performance up to 11.8%.
What’s more, The RCoV and RA scores for all three variants also revealed that even when faced with different perturbations, the model will demonstrate similar robustness performance, as shown in Table 3. This suggests that the robustness ability may be a common limitation for current LLMs, no matter what type the perturbation is.
6 Influential Factors for Believability
@@
We have already assessed the general believability performance of 10 LLMs. This section delves deeper into the four factors that exert substantial influences on believability. We anticipate that our studies could expedite subsequent research on human behavior simulation.
Simulation hallucination
As shown in Table 2, models demonstrate severe simulation hallucination with CA of Unknown questions is considerably lower than that of Known questions. One plausible possible explanation is that the model might have known the answer to a question due to the knowledge learned in the training process, even if the answer can not be deduced from the profile. Consequently, the model refuses to answer the question with "I do not know." as required in the prompt  [^6] In Appendix D, we further examine the effect of simulation hallucination by replacing the name of the character to compare the variants’ CA scores of Unknown questions.. This phenomenon reflects that models occasionally prefer to refuse or ignore the user’s instructions, which will greatly harm the user’s believability towards the model, especially when commercial simulation products are gaining increasing popularity, such as character.ai and npc.baichuan-ai.
Bias of models towards specific demographic traits
@@
We have found that believability can be significantly influenced by the profile perturbation in Section 5.3. Hence, it is a crucial inquiry: which profile information would yield high believability for various LLMs? To investigate this question, we compare the LLMs’ consistency by perturbing different demographic information in the profile. Specifically, we employ LLMs to simulate Homer by prompting the profile of Homer’s variants in the character variants dataset, whose profile is modified with only one demographic factor, such as birth year, while keeping all others unaltered.
@@
Table 4 shows the results. All LLMs exhibit various degrees of preference toward profiles with specific demographic factors. For instance, LongChat-7B-32K exhibits a significantly higher consistency score for the Caucasian race. Furthermore, we observe that profiles containing certain demographic factors correlate with higher consistency performance across a majority of models. Five models in the age group demonstrate superior performance when the birth year is 2000. Eight models in the race group demonstrate superior performance when the race is perturbed to the Caucasian compared to the other races. Within the group of surnames and educational backgrounds, models with the surname Bedonie and those who obtained Bachelor’s degrees exhibited similar higher performance. This observation indicates that different models consistently prefer specific demographic factors. This phenomenon may be attributed to the fact that many models are trained on overlapping corpora, resulting in the corpus bias being simultaneously manifested in all these models.
Position in the profile
@@
Model Known Unknown Normal Reverse Normal Reverse GPT-4 1.00 0.95 0.47 0.47 GPT-3.5-Turbo-16K 0.82 0.77 0.58 0.63 ChatGLM2-6B-32K 0.68 0.73 0.21 0.32 XVERSE-13B-Chat 0.68 0.73 0.53 0.53 Qwen-7B-Chat 0.64 0.64 0.11 0.11 Vicuna-13B-16K 0.64 0.68 0.32 0.37 Qwen-14B-Chat 0.59 0.73 0.21 0.26 LongChat-7B-32K 0.59 0.77 0.05 0.26 ChatGLM2-6B 0.50 0.59 0.16 0.32 Vicuna-7B-16K 0.36 0.64 0.05 0.11
@@
Table 5: The accuracy of Immutable Characteristic questions for models to simulate a character with the profile’s information order reversed (denoted as Reverse) and unchanged (denoted as Normal).
@@
For long textual inputs, models can pay different attention to the information in different positions. Hence, the believability is impacted by the placement of information inside the profile. To investigate this issue, we conduct experiments by adjusting the order of information in the profile. The original profile presents information in the order of Immutable Characteristic, Social Role, and Relationship, indicated as Normal. The adjusted order, denoted as Reverse, is Social Role, Relationship, and Immutable Characteristic. Then, we evaluate LLMs through the consistency dataset.
@@
Table 5 shows the results. The revised sequence order has significantly improved the CA scores of open-source models on the Immutable Characteristic questions. Nevertheless, this effect is not apparent for the commercial models. A possible explanation is that open-source models may struggle to adequately process lengthy textual content, even when their context size is large enough. Consequently, the model will allocate different attention to the information in the prompt’s different positions. Nevertheless, the commercial models retain strong processing capabilities when it comes to handling lengthy texts. Therefore, altering the order of sequence is less likely to influence the commercial model’s performance significantly.
Reasoning prompting
@@
Model Few Few+ CoT Few+ Self-Ask Zero Zero+ CoT GPT-4 0.77 0.77 0.82 0.75 0.77 GPT-3.5-Turbo-16K 0.70 0.77 0.77 0.77 0.77 XVERSE-13B-Chat 0.62 0.42 0.43 0.60 0.58 Qwen-14B-Chat 0.60 0.73 0.64 0.70 0.68 Vicuna-13B-16K 0.59 0.61 0.63 0.65 0.65 ChatGLM2-6B-32K 0.55 0.63 0.59 0.58 0.58 Qwen-7B-Chat 0.53 0.55 0.56 0.56 0.54 ChatGLM2-6B 0.49 0.54 0.49 0.44 0.41 LongChat-7B-32K 0.48 0.46 0.56 0.49 0.46 Vicuna-7B-16K 0.46 0.54 0.56 0.58 0.58
@@
Table 6: : The CA scores of models when simulating Homer with five different prompting strategies.
@@
Although reasoning prompting techniques, such as chain-of-thought, are considered effective in some tasks, we find they can not always increase the believability of human behavior simulation. To provide evidence, we conduct the simulation using prompt combinations of Few, Few+CoT, Few+Self-Ask, Zero, and Zero+CoT.
@@
Table 6 shows the results. Among all the prompt combinations considered, it is seen that no prompt combination exhibits a consistent improvement in the performance of all the models when compared to other prompts. One plausible explanation posits that the efficacy of these prompt techniques, such as CoT and Self-Ask, mostly lies in their ability to enhance performance on tasks involving reasoning abilities, such as solving, decision-making, and planning (Huang and Chang, 2022; Wang et al., 2022). Nevertheless, simulating human behaviors necessitates the model to hold other abilities, such as comprehensive comprehension of the character’s profile and the dynamics of character relationships.
@@
We also find that some open-source models, such as the Qwen-14B-Chat and the Vicuna series, perform even better when no demonstration examples are included in the prompt (Zero) compared with the Few setting. We carefully analyzed their responses and found that these models consistently generate the exemplars in the Few setting as a response. One potential reason is that the lengthy profile and the challenging task complexity hinder the model from comprehending the exemplar in the Few setting.